solve the differential equation dy/dx + (x/1-x^2)y=xy^1/2 , y(0)=1

correct question
1
Expert's answer
2020-05-19T08:56:48-0400

the given equation can be written as

"\frac{1}{\sqrt{y}} \frac{dy}{dx} + \frac{2x\sqrt{y}}{2(1-x^2)} = x" .....(1)


this is of the form f'(y) y' + P f(y) = Q

where P = "\frac{x}{2(1-x^2)}"

Q= x

f(y) = 2 "\sqrt{y}"


let v= 2"\sqrt{y}"

differentiating,

"\frac{1}{\sqrt{y}} \frac{dy}{dx} = \frac{dv}{dx}"


(1) becomes

"\frac{dv}{dx} + \frac{xv}{2(1-x^2)} = x"


this is of the form "\frac{dv}{dx} + Av = B"


"A = \frac{x}{2(1-x^2)} and B=x"


integrating factor = "e^{\intop Pdx} = e^{\intop \frac{x}{2(1-x^2)}dx} = (1-x^2)^{(-1/4)}"


the general solution is given by:

v (1-x2)-1/4 = "\int" x. (1-x2)-1/4 dx + c


2"\sqrt{y}" = -2/3 (1-x2) + c(1-x2)1/4

now, given y(0)=1


2 = -2/3 + c

c= 8/3


so, the solution is

2"\sqrt{y}" = -2/3 (1-x2) + 8/3(1-x2)1/4